# ShanghaiTech Mapping Robot is All You Need: Robot System for Collecting Universal Ground Vehicle Datasets

### v1

@@ III-G Ultrasound Range Finders
Ultrasound Range Finders are often equipped by road vehicles for detecting near-field obstacles especially in the context of automatic parking. We equipped our robot with 13 KS103 ultrasonic range finders. KS103 has a range of 1cm to 8m and a resolution of 1cm. It utilizes proprietary filtering techniques to improve ranging accuracy, claiming high precision even in noisy environments.
Figure 6: MTi-630R IMU (left) and KS103 ultrasound range finder (right)
@@
Figure 7: TI IWR6843AOPEVM mmWave Radar
For our robot, we chose IWR6843AOPEVM, a 60GHz mmWave radar from TI. It achieves a wide 120° azimuth and 120° elevation field of view with a small antenna-on-package module. We decided to have five of these radars to look at front, back, left, right, and up directions.
@@ III-I GNSS receiver
We employed a Bynav A1 GNSS RTK receiver that utilizes GPS, GLONASS, Galileo, BeiDou, QZSS, and NAVIC constellations. The receiver supports dual antenna setup, measuring not only the position but also the attitude of the robot. It also has an builtin INS that helps improve positioning both in real-time and for processing afterwards.
@@ V System integration & Sensor Setup
Figure 13: A simple model of ShanghaiTech Mapping Robot II’s mechanical design, showing Clearpath Husky UGV, the sensor deck, front sensor tower and back sensor tower.
@@ V-A Front Sensor Tower
As showed in Fig. 15 and Fig. 20, PandarQT deserves a position on the very top to observe robot’s near surrounding, including objects at very low angle, with no obstruction. Stayed straight under PandarQT, the front Ruby covers far away objects. Both LiDARs, as well as all rest of them, are connected to separate cluster nodes through Ethernet cable. Each of these nodes handles the control and data collection of the respective LiDAR. Robosense LiDARs are synced using the LiDAR synchronization channel mentioned in Sec. IV-C while LiDARs of other models use PTP protocol to get time base from the syncboard relayed by the cluster nodes.
@@
Figure 15: The front sensor tower of ShanghaiTech Mapping Robot II.
@@
KS103 ultrasound range finders are installed on curved racks attached to the tower to achieve evenly angled detecting directions. All 13 of them are connected via an I2C bus to a microcontroller which triggers them in an interleaved sequence to avoid cross-talk.
@@
At the front of the UGV base, there is a downward looking Grasshopper3 camera, as showed in Fig. 16. Its focal distance is set to match the distance to the ground in order to get a clear view of the floor, road, etc. Alongside it, two 1.5-watt LED lights are installed to ensure enough lighting, enabling a even higher frame rate to compensate for the limited FoV (about $40\times 60$ cm on the ground) due to its low mounting position.
@@ V-B Sensor Deck
Figure 17: The design of ShanghaiTech Mapping Robot II’s sensor deck. Mounting position for Grasshopper3 (blue), Azure Kinect DK (green), computation cluster (yellow), front sensor tower (purple) and back sensor tower (red).
@@
As showed in Fig. 18, each mmWave radar is mounted on the back of a Azure Kinect DK using a 3D printed part, forming a Kinect-mmWave pair. Such pairs are attached to Kinect mounting slots of sensor deck, marked green in Fig. 17. Such approach insures a uniform translation between all 5 pairs thus helps simplifying the calibration of mmWave radar extrinsics.
@@ V-C Back Sensor Tower
Figure 19: The back sensor tower of ShanghaiTech Mapping Robot II viewed from its back (left) and right (right).
@@
Figure 20: FoV of LiDARs equipped by ShanghaiTech Mapping Robot II.
@@
Figure 21: The communication structure of the ShanghaiTech Mapping Robot II.
Figure 22: The synchronization flow of the ShanghaiTech Mapping Robot II.
@@ VI Calibration
To ensure accurate perception and mapping, it is crucial to calibrate the intrinsic and extrinsic parameters of the various sensors employed in the system. This section outlines the calibration procedures for the RGB cameras, event-based cameras, and LiDAR sensors.
@@ VI-B LiDAR Extrinsic Calibration
To calibrate the extrinsics between LiDARs, we registered the scans of each LiDAR against a high-precision scan from a FARO® Focus 3D scanner. The LiDAR scans were captured with the robot standing still and then averaged to reduce temporal noise. The registration algorithm, Iterative Closest Point (ICP) to be specific, yields the transformations between LiDARs. Combined with the camera-LiDAR transformation, all LiDARs were connected to the extrinsics tree.
@@ VII Procedures and Post-Processing
The mapping robot is designed with distributed computation in mind so it is vital to orchestrate all the nodes in the computation cluster along with the sensors connected to them. In this section we would introduce the procedure of data collection and post-process to be performed on the collected data.
VII-A Procedure of data collection
@@ VII-A1 Material Preparation
At minimum, the robot itself, one battery, and a controlling terminal like KVM or laptop is needed for performing a collection run. A second battery as well as backup ones or power adaptor is recommended.
@@ VII-A5 Sensors Bringup
The master node orchestrates all worker nodes to launch the sensor drivers. We would manually check if all sensors work as intended. If that’s not the case, sensor’s parameter would be modified accordingly. After everything is confirmed to work well, the triggering is turned off await the upcoming data collection.
VII-A6 Data collection
Command the master node to orchestrate a distributed recording of sensor data. Turn on triggering to officially start a data collection run. We would now drive the robot along the data collection route.
VII-A7 End of collection
After the robot reaches finishing position of the route, the triggering is turned off first, followed by the data recording processes. Sensor drivers stay running and do not have to be spawned repetitively until there is no more data collection runs to be carried out before powering off the robot. The data collected can be copied either into external hard drives via USB port from each nodes or to a workstation or network storage device through Ethernet connection, pending for post-processing.
@@ VII-B3 Image Compression
To provide an alternative compressed version for faster distribution of our dataset, we encode each image stream using AV1 codec.
@@ VII-B4 Mapping Ground Truth
A FARO® Focus 3D scanner is used to obtain a precise point cloud of the environment along the data collection route, acting as the mapping or 3D reconstruction ground truth.
@@ VII-B5 Trajectory Ground Truth
VIII Experiments
VIII-A LiDAR&RGBD Interference Evaluation
@@
VIII-A1 Precision analysis
@@
TABLE I: Mean & S.D. of temporal S.D. of distance (m)
@@
Table I shows that all but the two Robosense Bpearls and the Livox Avia have consistent precision between the sole and collective runs, indicating that the precision of the other LiDARs is not affected by interference.
@@
Fig. 26 (a) & (b) shows the temporal mean of the point cloud collected by the front Robosense Bpearl. In the collective run, both Robosense Bpearl and Livox Avia give noisy points behind the wall or below the floor, indicating interference. However, the non-noisy points have consistent standard deviation of distance compared to the sole run, suggesting that with proper filtering, the two sensors would give point clouds of the same precision regardless of interference.
VIII-A2 Accuracy analysis
@@
TABLE II: Mean & S.D. of Point-to-Plane distance (m)
@@
Table II shows that the difference in point-to-plane distance between the sole and collective runs lies within $2\sigma$ for all LiDARs, indicating no significant accuracy change due to interference.
TABLE III: Pixel-wise Mean & S.D. of Distance Difference(m)
@@
For depth cameras, the difference in average distance measurements between the sole and collective runs is calculated pixel-wise to detect accuracy changes. As showed in Table III, the mean difference is much smaller than its standard deviation, indicating that the depth cameras’ accuracy is also unaffected by interference.
@@
VIII-B Demo Dataset
To assess the mapping robot’s data collection capability and verify the compatibility of the collected data with SLAM algorithms, we conducted an outdoor experiment on the ShanghaiTech University campus. The experiment involved collecting data along a predefined route and running SLAM algorithms on the collected data.
VIII-B1 Data Collection
Figure 27: Route plan for outdoor data collection experiment
The 1.4 km data collection route, shown in Fig. 27, was designed to include a variety of terrains and scenes on the ShanghaiTech University campus. The route was evenly split between areas around buildings and parkland, with the mapping robot primarily driving on paved roads or solid structures like bridges, and briefly on grass. Three sections of the route, surrounded by buildings, woods, or a blend of both, were revisited after traveling elsewhere, providing loop-closing opportunities for SLAM algorithms. The route also included slopes to enrich the elevation diversity.
Figure 28: Mapping robot performing data collection on ShanghaiTech campus
The experiment was conducted on an overcast day, providing even lighting on objects and preventing direct sunlight spots in the collected images. The robot took 26 minutes to traverse the entire route, ultimately collecting 2,618 GiB of data, as shown in Fig. 29. A detailed list of the recorded ROS topics is provided in the Appendix.
Figure 29: Ground truth trajectory of the experiment run, post-processed from differential GNSS and IMU data
VIII-B2 Lidar-based SLAM
VIII-B3 Visual SLAM
IX Conclusions

### v2

@@ III-G Ultrasound Range Finders
Ultrasound Range Finders are often equipped by road vehicles for detecting near-field obstacles especially in the context of automatic parking. We equipped our robot with 13 KS103 ultrasonic range finders (Fig. 2-m) . KS103 has a range of 1cm to 8m and a resolution of 1cm. It utilizes proprietary filtering techniques to improve ranging accuracy, claiming high precision even in noisy environments.
@@
For our robot, we chose IWR6843AOPEVM (Fig. 2-n) , a 60GHz mmWave radar from TI. It achieves a wide 120° azimuth and 120° elevation field of view with a small antenna-on-package module [40]. We decided to have five of these radars to look at front, back, left, right, and up directions.
@@ III-I GNSS receiver
We employed a Bynav A1 GNSS RTK receiver (Fig. 2-f) that utilizes GPS, GLONASS, Galileo, BeiDou, QZSS, and NAVIC constellations. The receiver supports dual antenna setup, measuring not only the position but also the attitude of the robot. It also has an builtin INS that helps improve positioning both in real-time and for processing afterwards.
@@ V System Integration & Sensor Setup
Figure 8: A simplified model of ShanghaiTech Mapping Robot ’s mechanical design, showing Clearpath Husky UGV, the sensor deck, front sensor tower and back sensor tower.
@@ V-A Front Sensor Tower
As showed in Fig. 10 and Fig. 15, PandarQT deserves a position on the very top to observe robot’s near surrounding, including objects at very low angle, with no obstruction. Stayed straight under PandarQT, the front Ruby covers far away objects. Both LiDARs, as well as all rest of them, are connected to separate cluster nodes through Ethernet cable. Each of these nodes handles the control and data acquisition of the respective LiDAR. Robosense LiDARs are synced using the LiDAR synchronization channel mentioned in Sec. IV-C while LiDARs of other models use PTP protocol to get time base from the syncboard relayed by the cluster nodes.
@@
Figure 10: The front sensor tower of ShanghaiTech Mapping Robot.
@@
MTi-630R is configured to measure linear acceleration, angular velocity, orientation as well as magnetic heading, atmospheric pressure and temperature, all at a rate of up to 400Hz triggered by Syncboard. KS103 ultrasound range finders are installed on curved racks attached to the tower to achieve evenly angled detecting directions. All 13 of them are connected via an I2C bus to a microcontroller which triggers them in an interleaved sequence to avoid cross-talk.
@@
At the front of the UGV base, there is a downward looking Grasshopper3 camera, as showed in Fig. 11. Its focal distance is set to match the distance to the ground in order to get a clear view of the floor, road, etc. Alongside it, two 1.5-watt LED lights are installed to ensure enough lighting, enabling a even higher frame rate to compensate for the limited FoV (about $40\times 60$ cm on the ground) due to its low mounting position. We dedicated a triggering channel of Syncboard for this camera to drive it at a different, often faster, frame rate than its counterparts looking around and upward.
@@ V-B Sensor Deck
Figure 12: The design of ShanghaiTech Mapping Robot’s sensor deck. Mounting position for Grasshopper3 (blue), Azure Kinect DK (green), computation cluster (yellow), front sensor tower (purple) and back sensor tower (red).
@@
As showed in Fig. 13, each mmWave radar is mounted on the back of a Azure Kinect DK using a 3D printed part, forming a Kinect-mmWave pair. Such pairs are attached to Kinect mounting slots of sensor deck, marked green in Fig. 12. Such approach insures a uniform translation between all 5 pairs thus helps simplifying the calibration of mmWave radar extrinsics. While mmWave radars do not support hardware triggering, Azure Kinect DKs do. They are all configured to run in the subordinate mode and have their sync-in ports connected to a dedicated triggering channel of Syncboard generating VSync signal at desired frame rate.
@@ V-C Back Sensor Tower
Figure 14: The back sensor tower of ShanghaiTech Mapping Robot viewed from its back (left) and right (right).
@@
Figure 15: FoV of LiDARs equipped by ShanghaiTech Mapping Robot.
@@
Figure 16: The communication structure of the ShanghaiTech Mapping Robot.
Figure 17: The synchronization flow of the ShanghaiTech Mapping Robot.
@@ VI Calibration
To ensure accurate perception and mapping, it is crucial to calibrate the intrinsic and extrinsic parameters of the various sensors employed in the system. There are various tools suitable for our use case, including ones concentrate on sensors of one same type [42], [43], [44], and others performing cross-modal calibration [45], [46], [47] and [48]. This section outlines the calibration procedures for the RGB cameras, event-based cameras, and LiDAR sensors.
@@ VI-B LiDAR Extrinsic Calibration
To calibrate the extrinsics between LiDARs, we registered the scans of each LiDAR against a high-precision scan from a FARO Focus 3D scanner [50]. The LiDAR scans were captured with the robot standing still and then averaged to reduce temporal noise. The registration algorithm, Iterative Closest Point (ICP) to be specific, yields the transformations between LiDARs. Combined with the camera-LiDAR transformation, all LiDARs were connected to the extrinsics tree.
@@ VII Procedures and Post-Processing
The mapping robot is designed with distributed computation in mind so it is vital to orchestrate all the nodes in the computation cluster along with the sensors connected to them. In this section we would introduce the procedure of data acquisition and post-process to be performed on the collected data.
VII-A Procedure of data acquisition
@@ VII-A1 Material Preparation
At minimum, the robot itself, one battery, and a controlling terminal like KVM or laptop is needed for performing a acquisition run. A second battery as well as backup ones or power adaptor is recommended.
@@ VII-A5 Sensors Bringup
The master node orchestrates all worker nodes to launch the sensor drivers. We would manually check if all sensors work as intended. If that’s not the case, sensor’s parameter would be modified accordingly. After everything is confirmed to work well, the triggering is turned off await the upcoming data acquisition.
VII-A6 Data acquisition
Command the master node to orchestrate a distributed recording of sensor data. Turn on triggering to officially start a data acquisition run. We would now drive the robot along the data acquisition route.
VII-A7 End of acquisition
After the robot reaches finishing position of the route, the triggering is turned off first, followed by the data recording processes. Sensor drivers stay running and do not have to be spawned repetitively until there is no more data acquisition runs to be carried out before powering off the robot. The data collected can be copied either into external hard drives via USB port from each nodes or to a workstation or network storage device through Ethernet connection, pending for post-processing.
@@ VII-B3 Image Compression
To provide an alternative compressed version for faster distribution of our dataset, we encode each image stream using AV1 codec [52].
@@ VII-B4 Mapping Ground Truth
A FARO Focus 3D scanner is used to obtain a precise point cloud of the environment along the data acquisition route, acting as the mapping or 3D reconstruction ground truth.
@@ VII-B5 Trajectory Ground Truth
VIII LiDAR&RGBD Interference Evaluation
@@
VIII-A Precision analysis
@@
TABLE II: Mean & S.D. of temporal S.D. of distance (m)
@@
Table II shows that all but the two Robosense Bpearls and the Livox Avia have consistent precision between the sole and collective runs, indicating that the precision of the other LiDARs is not affected by interference.
@@
Fig. 2121(a) & 21(b) shows the temporal mean of the point cloud collected by the front Robosense Bpearl. In the collective run, both Robosense Bpearl and Livox Avia give noisy points behind the wall or below the floor, indicating interference. However, the non-noisy points have consistent standard deviation of distance compared to the sole run, suggesting that with proper filtering, the two sensors would give point clouds of the same precision regardless of interference.
VIII-B Accuracy analysis
@@
TABLE III: Mean & S.D. of Point-to-Plane distance (m)
@@
Table III shows that the difference in point-to-plane distance between the sole and collective runs lies within $2\sigma$ for all LiDARs, indicating no significant accuracy change due to interference.
TABLE IV: Pixel-wise Mean & S.D. of Distance Difference(m)
@@
For depth cameras, the difference in average distance measurements between the sole and collective runs is calculated pixel-wise to detect accuracy changes. As showed in Table IV, the mean difference is much smaller than its standard deviation, indicating that the depth cameras’ accuracy is also unaffected by interference.
@@
IX Outdoor Dataset
To assess the mapping robot’s data acquisition capability and verify the compatibility of the collected data with SLAM algorithms, we conducted an outdoor experiment on the ShanghaiTech University campus. The experiment involved collecting data along a predefined route and running SLAM algorithms on the collected data.
IX-A Data Acquisition
An 1.4 km data acquisition route, as shown in Fig. 23, was designed to include a variety of terrains and scenes on the ShanghaiTech University campus. The route was evenly split between areas around buildings and parkland, with the mapping robot primarily driving on paved roads or solid structures like bridges, and briefly on grass. Three sections of the route, surrounded by buildings, woods, or a blend of both, were revisited after traveling elsewhere, providing loop-closing opportunities for SLAM algorithms. The route also included slopes to enrich the elevation diversity.
Figure 22: Mapping robot performing data acquisition on ShanghaiTech campus
The experiment was conducted on an overcast day, providing even lighting on objects and preventing direct sunlight spots in the collected images. The robot took 26 minutes to traverse the entire route, ultimately collecting 2,618 GiB of data, as shown in Fig. 23. A detailed list of the recorded ROS topics is provided in the Appendix A.
Figure 23: Ground truth trajectory of the outdoor dataset, with its start (green) and end (red) point marked out. Color of the trajectory goes from light yellow to dark orange as time progresses.
IX-B Lidar-based SLAM
IX-C Visual SLAM
X Conclusions
XI Acknowledgments
This work was supported by Science and Technology Commission of Shanghai Municipality (STCSM), project 22JC1410700 “Evaluation of real-time localization and mapping algorithms for intelligent robots”. This work has also been partially funded by the Shanghai Frontiers Science Center of Human-centered Artificial Intelligence.
@@ Appendix A Details of ROS topics recorded in experiment
In Table. V, we showed for each type and model of sensor the message topics they publish to as well as the total number of messages recorded from them. The time duration, data rate, and message type are also listed.
TABLE V: Sensor Data Statistics of Outdoor Dataset
Sensor Type Model Topic Topic Count Messages Data Rate Duration Message Type
[…]
FLIR LadyBug5+ /ladybug/camera{0-5}/bayer{0-3}/compressed 24 23,430 15 Hz 26:01 sensor_msgs/CompressedImage
[…]
Hesai PandarQT /hesai/pandar 1 16,031 10 Hz 26:41 sensor_msgs/PointCloud2
Solid-state LiDAR Livox Avia /livox/lidar 1 16,038 10 Hz 26:43 sensor_msgs/PointCloud2
/livox/imu 1 325,608 200 Hz 26:43 sensor_msgs/Imu
Neuvition M1 /neuvition_cloud 1 1,489 10 Hz 2:27 sensor_msgs/PointCloud2
/neuvition_image/compressed 1 859 5 Hz 2:27 sensor_msgs/CompressedImage
IMU Xsens MTi-630R /mti/imu/data 1 624,797 400 Hz 26:02 sensor_msgs/Imu
/mti/filter/free_acceleration 1 624,797 400 Hz 26:02 geometry_msgs/Vector3Stamped
/mti/filter/quaternion 1 624,797 400 Hz 26:02 geometry_msgs/QuaternionStamped
/mti/imu/acceleration 1 624,797 400 Hz 26:02 geometry_msgs/Vector3Stamped
/mti/imu/angular_velocity 1 624,797 400 Hz 26:02 geometry_msgs/Vector3Stamped
/mti/imu/dq 1 624,797 400 Hz 26:02 geometry_msgs/QuaternionStamped
/mti/imu/dv 1 624,797 400 Hz 26:02 geometry_msgs/Vector3Stamped
/mti/imu/mag 1 624,797 400 Hz 26:02 geometry_msgs/Vector3Stamped
/mti/imu/time_ref 1 624,797 400 Hz 26:02 sensor_msgs/TimeReference
/mti/pressure 1 624,797 400 Hz 26:02 sensor_msgs/FluidPressure
/mti/temperature 1 624,797 400 Hz 26:02 sensor_msgs/Temperature
Ultrasonic Sensor KS103 /ultrasound/range/{1-13} 13 1,387 1 Hz 26:47 sensor_msgs/Range
[…]
GNSS Receiver Bynav A1 /gps/fix 1 8,045 5 Hz 26:49 sensor_msgs/NavSatFix
/gps/bestpos 1 8,045 5 Hz 26:49 novatel_gps_msgs/NovatelPosition
/gps/bestvel 1 8,045 5 Hz 26:49 novatel_gps_msgs/NovatelVelocity
/gps/gpgga 1 8,045 5 Hz 26:49 novatel_gps_msgs/Gpgga
/gps/gprmc 1 8,045 5 Hz 26:49 novatel_gps_msgs/Gprmc
/gps/gps 1 7,988 5 Hz 26:49 gps_common/GPSFix
/gps/inspvax 1 8,046 5 Hz 26:49 novatel_gps_msgs/Inspvax
Synchronization Syncboard /syncboard/line/{2,14-17} 5 7,810-624,786 5-400 Hz 26:40 std_msgs/Time
/syncboard/status 1 1,601 1 Hz 26:40 mars_syncboard_srvs/BoardStatus
Robot Base Clearpath Husky /joy_teleop/joy_selected 1 37,980 24 Hz 26:41 sensor_msgs/Joy
/status 1 1,602 1 Hz 26:41 husky_msgs/HuskyStatus
Total 137 # of bags: 37, size of bags: 2,618 GiB